If log 4 (log 2 (logx)) = ½ then x = ? a) 9 b)27 c)81 d)243

If log4(log2(logx)) = ½ then x = ?
a) 9
b)27
c)81
d)243

Grade:9

2 Answers

Pushparaj Patel
8 years ago
If the base of the log is 3 than the ans is opt. C) 81
Okk
 
log4(log2(log3x))=1/2
log2(log3x)=41/2=2
log3x=22=4
x=34=81
 
Have a Nice Day...
Pushparaj Patel
8 years ago
Whether the ans is right or not??? Just varify from respective teache okk